# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 3. Some explicit polarizations and skeleta
We start our computation of explicit polarizations with the simplest case, in dimension $2$. This case is then used in §$\S$ 3.2 to construct Liouville polarizations on bidiscs. In §$\S$ 3.3 we show how certain singular Liouville polarizations can be surgered to smooth Liouville polarizations.
3.1. Polarizations in dimension $2$
@@
Let $(S,\omega)$ be a 2-dimensional compact symplectic surface with non-empty boundary $\partial S$. We define a Liouville polarization of $S$ to be a finite set of points $p_{i}$ in $\operatorname{Int}S$ with weights $\mu_{i}>0$ together with a tame Liouville form $\lambda$ defined on $S\setminus\bigcup_{i}p_{i}$ such that $X_{\lambda}$ is tangent to $\partial S$. Tameness now means that for each $i$ there are symplectic polar coordinates $(R,\theta)$ near $p_{i}$ such that $\lambda=-\mu_{i}d\theta+\lambda^{\prime}$ on $\operatorname{Op}(p_{i})\setminus\{p_{i}\}$, where $\lambda^{\prime}$ is a bounded 1-form on $\operatorname{Op}(p_{i})\setminus\{p_{i}\}$. One easily checks that $\sum_{i}\mu_{i}={\mathcal{A}}_{\omega}(S)$. Conversely, any finite set of positively weighted points $\{(p_{i},\mu_{i})\}$ on $(S,\omega)$ that satisfies $\sum_{i}\mu_{i}={\mathcal{A}}_{\omega}(S)$ admits a Liouville polarization, as is easy to see. Because of this very direct link between the residues and the area in dimension $2$, we switch notation from $\mu_{i}$ to $a_{i}$ throughout this section. Its main purpose is to explain that in dimension $2$, the skeleton can be prescribed, which seems much harder in higher dimensions. We recall the notion of regular grid in a slightly more general setting.
@@
Let $(S,\omega)$ be a compact surface with boundary, and let $\Gamma$ be a regular grid that decomposes $S$ into $m$ discs of area $a_{i}$, in each of which we chose a point $p_{i}$. Then there exists a Liouville polarization ($\bigl{(}\{(p_{i},a_{i})\},\lambda\bigr{)}$ of $S$ whose skeleton is $\Gamma$. This Liouville form can be required to coincide with $(R-a_{i})d\theta$ in Darboux coordinates near each $p_{i}$.
In the case of a closed disc $\overline{D}(A)$ (which will be the example we will use later on) one may attempt to prove the proposition as follows: Just define a first tame Liouville form with residues $-a_{i}$ at $p_{i}$ by $\lambda:=\alpha_{{\operatorname{st}}}-\sum a_{i}\,d\theta_{i}$, where $\theta_{i}$ is the pull-back by the translation of the vector $-p_{i}$ of the standard angular coordinate in $\mathbb{C}\setminus\{0\}$. Then the periods of $\lambda$ vanish on each loop of $\Gamma$, so we can correct $\lambda$ by adding the differential of a smooth function in order to make $\lambda$ vanish on $\Gamma$. Then the Liouville flow of $\lambda$ fixes $\Gamma$ pointwise, and so ($\bigl{(}\{(p_{i},a_{i})\},\lambda\bigr{)}$ provides a Liouville polarization of $\overline{D}(A)$ whose skeleton contains $\Gamma$. Unfortunately, this proof does not constrain the skeleton enough: it could be strictly bigger than $\Gamma$. We therefore proceed in a different manner. Recall that $X_{\lambda}$ is tangent to $\ker\lambda$, since $\lambda(X_{\lambda})=d\lambda(X_{\lambda},X_{\lambda})=0$. Our strategy is therefore to first construct a (singular) foliation ${\mathcal{F}}$ to which we wish $X_{\lambda}$ to be tangent to. Once this “vanishing foliation” is constructed, the following lemma will readily provide the Liouville form itself.
@@
for the set of singularities of $\Gamma$ in the boundary of ${\mathcal{D}}_{i}$. For each $p_{i}$ and a sufficiently small $\varepsilon$ the translation $(D(p_{i}),p_{i})\to(D(\varepsilon),0)$ is a Darboux chart that induces coordinates $(R,\theta)$ on $D(p_{i})$. Similarly, by our assumption on $\Gamma$ we can consider for each $q\in\operatorname{Sing}(\Gamma)\setminus\partial D$ symplectic coordinates $\phi_{q}=(R,\theta)\colon D(q)\to D(\varepsilon)$ on a disk around $q$ such that
@@
Also, when $q\in\operatorname{Sing}(\Gamma)\cap\partial D$, we have symplectic coordinates $(R,\theta)$ on a neighbourhood $D^{+}(q)$ of $q$ in $D$, with values in $\{R<\varepsilon,\theta\in[0,\pi]\}$ such that
@@
The radial ray that cuts the sector $S_{i}(q)$ into two equal sectors is called a local separatrix at $q$ and is denoted $s_{i}(q)$.
@@
For a point $q\in\operatorname{Sing}(\Gamma)\setminus\partial D$, let $\rho_{q}\colon D(\varepsilon)\to\mathbb{C}$ be the “ramified covering”
@@
Note that $\rho_{q}$ is smooth and symplectic on $D(\varepsilon)\setminus\{0\}$, but only continuous at 00. The sectors $S_{i}(q)$ around $q$ are the preimages under $\rho_{q}\circ\phi_{q}$ of $\{y\geq 0\}$ or $\{y\leq 0\}$, and the separatrices $s_{i}(q)$ are sent by $\rho_{q}\circ\phi_{q}$ to $\{x=0\}$. We define the foliation in $D(q)$ to be the pull-back by $\rho_{q}\circ\phi_{q}$ of the vertical foliation:
@@
    As a result, the foliation ${\mathcal{F}}$ can be extended to a smooth foliation on $\textnormal{Op}\,(\Gamma)\backslash\operatorname{Sing}(\Gamma)$ that is transverse to the regular part of $\Gamma$.
@@
Figure 3.1. Notation, and the foliation on $D(\varepsilon)$ for $m_{q_{1}}=3$.
@@
    On the $D(p_{i})$, $D(q_{j})$, $D^{+}(q_{k})$, and $\textnormal{Op}\,(\Gamma)\backslash\operatorname{Sing}(\Gamma)$ the leaves of ${\mathcal{F}}$ are the ones described in Step 1.
@@
Since ${\mathcal{F}}_{i}$ is a smooth foliation on ${\mathcal{D}}_{i}\setminus\bigl{(}\{p_{i}\}\cup{\mathcal{S}}_{i}\bigr{)}$ and by the smooth fitting of the leaves across the $\partial{\mathcal{D}}_{i}(q_{j}\to q_{j+1})$, the foliation ${\mathcal{F}}$ on $D\setminus\bigl{(}\cup\{p_{i}\}\bigcup\operatorname{Sing}(\Gamma)\bigr{)}$ defined by
@@
is $C^{\infty}$-smooth, and has a unique extension to a singular foliation on the whole disc $D$ whose singularities are $\cup\{p_{i}\}\bigcup\operatorname{Sing}(\Gamma)$. The different separatrices $s(p_{i}\to q)$, $q\in{\mathcal{S}}_{i}$, are smooth leaves. The leaves of ${\mathcal{F}}$ that pass through a point in $\Gamma\setminus\operatorname{Sing}(\Gamma)$ that separates between ${\mathcal{D}}_{i}$ and ${\mathcal{D}}_{j}$ join $p_{i}$ to $p_{j}$. Finally, the leaves through a sequence of regular points on ${\mathcal{D}}_{i}\cap{\mathcal{D}}_{j}$ that converge to a singular point $q\in\operatorname{Sing}(\Gamma)$ break into the two separatrices $s(p_{i}\to q)$ and $s(p_{j}\to q)$.
@@
[…]
@@
[…]
$\lambda|_{\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})}\,:=\,\lambda_{i}=\lambda_{j}.$
[…]
@@
[…]
@@
    $\lambda$ is tame at each $p_{i}$ with residue $-a_{i}$, in fact $\lambda=(R-a_{i})d\theta$ near $p_{i}$,
@@
    $X_{\lambda}$ vanishes on $\Gamma\backslash\operatorname{Sing}\Gamma$,
@@
We only treat the case of $q\in\operatorname{Sing}(\Gamma)\setminus\partial D$, the other case being similar. We recall that a neighbourhood $D(q)$ of $q$, that is identified with $D(\varepsilon)$, covers a neighbourhood of $0\in\mathbb{C}$ via the map $\rho_{q}(R,\theta)=(\frac{2R}{m_{q}},\frac{m_{q}}{2}\theta)$, and that in these coordinates, ${\mathcal{F}}$ is the foliation $\rho_{q}^{*}\{x=\text{const}\}$. This foliation is thus also tangent to the kernel of the local Liouville form $\lambda_{q}:=\rho_{q}^{*}(-\tfrac{1}{\pi}\;\!y\>\!dx)$. Away from $q$, this form $\lambda_{q}$ also vanishes on $\rho_{q}^{*}\{y=0\}=\Gamma\cap D(q)$. Away from $q$, both forms $\lambda$ and $\lambda_{q}$ are thus smooth and vanish along ${\mathcal{F}}$ and on $\Gamma$. Stokes’ theorem therefore guarantees that they coincide on $D(q)\backslash\{q\}$. None of the two forms smoothly extends to $q$, however, since at $(x,y)\neq(0,0)$,
@@
The function $r^{m_{q}}\sin(2\pi m_{q}\theta)$ is smooth on $\mathbb{R}^{2}$, since it is the imaginary part of $z\mapsto z^{m_{q}}$. Let $\chi\colon[0,2\varepsilon]\to\mathbb{R}_{\geq 0}$ be a function that coincides with $r^{m_{q}}$ near 00 and with $R$ on $[\varepsilon,2\varepsilon$], is positive and smooth except at 00, and meets $\chi(R)<R$ on $(0,\varepsilon)$. Define the smooth 1-form $\lambda^{\prime}$ on $D$ by $\lambda^{\prime}=\lambda$ on $D\setminus\cup_{q\in\operatorname{Sing}\Gamma}D(q)$ and
@@
Then $\lambda^{\prime}$ is a smooth Liouville form on $D\setminus\{p_{1},\dots,p_{m}\}$, tame at $p_{i}$ with residue $-a_{i}$. We claim that the skeleton of $\lambda^{\prime}$ is $\Gamma$, as required. To see this, notice that $X_{\lambda^{\prime}}=X_{\lambda}$ on $D\setminus\cup_{q\in\operatorname{Sing}\Gamma}D(q)$, and
@@
$X_{\lambda^{\prime}}$ is radial on $\Gamma\cap D(q)$ (hence its flow preserves $\Gamma$), and the trajectories of the points in $D(q)\setminus\Gamma$ flow away from $\Gamma$. From the $\frac{\partial}{\partial R}$-component, and from $\Gamma\supset\{\cos(2\pi m_{q}\theta)=1\}$ and $\chi(R)<R$ on $(0,\varepsilon)$, we infer that the trajectories of the points in $D(q)\setminus\Gamma$ all leave $D(q)$, and hence reach the set $\{\lambda=\lambda^{\prime}\}\setminus\Gamma$ (since $\Gamma$ is invariant). It follows that $\textnormal{Skel}(\lambda^{\prime})=\Gamma$. $\square$
@@
Let $(S,\omega)$ be a compact symplectic surface with or without boundary. Let $\bigl{(}{\mathcal{P}}:=\{(p_{i},a_{i}),\;i=1,\dots,n\},\lambda\bigr{)}$ with $p_{i}\in S\setminus\partial S$ be a polarization of $S$ (where $\lambda$ is a tame Liouville form with residues $a_{i}$ at $p_{i}$ and with flow tangent to $\partial S$), and set $\Gamma:=\textnormal{Skel}(S,{\mathcal{P}},\lambda)$.
@@
Then $\lambda^{\prime}$ is a Liouville form tame at $p_{i}^{j}$ with residue $a_{i}^{j}$. Since it coincides with $\lambda$ on $\Gamma\cup\partial S$, its flow preserves both $\Gamma$ and $\partial S$, so $\left({\mathcal{P}}^{\prime}:=\left\{(p_{i}^{j},a_{i}^{j})\right\},\lambda^{\prime}\right)$ is a Liouville polarization of $S$. Since $\lambda^{\prime}$ coincides with $\lambda$ outside the discs $D_{i}$,

### v2

@@ 3. Some explicit polarisations and skeleta
We start our computation of explicit polarisations with the simplest case, in dimension $2$. This case is then used in $\S$ 3.2 to construct Liouville polarisations on bidiscs. In $\S$ 3.3 we show how certain singular Liouville polarisations can be surgered to smooth Liouville polarisations.
3.1. Polarisations in dimension 22
@@
Let $(S,\omega)$ be a 2-dimensional compact symplectic surface with non-empty boundary $\partial S$. We define a Liouville polarisation of $S$ to be a finite set of points $p_{i}$ in $\operatorname{Int}S$ with weights $\mu_{i}>0$ together with a tame Liouville form $\lambda$ defined on $S\setminus\bigcup_{i}p_{i}$ such that $X_{\lambda}$ is tangent to $\partial S$. Tameness now means that for each $i$ there are symplectic polar coordinates $(R,\theta)$ near $p_{i}$ such that $\lambda=-\mu_{i}d\theta+\lambda^{\prime}$ on $\operatorname{Op}(p_{i})\setminus\{p_{i}\}$, where $\lambda^{\prime}$ is a bounded 1-form on $\operatorname{Op}(p_{i})\setminus\{p_{i}\}$. One easily checks that $\sum_{i}\mu_{i}={\mathcal{A}}_{\omega}(S)$. Conversely, any finite set of positively weighted points $\{(p_{i},\mu_{i})\}$ on $(S,\omega)$ that satisfies $\sum_{i}\mu_{i}={\mathcal{A}}_{\omega}(S)$ admits a Liouville polarisation, as is easy to see. Because of this very direct link between the residues and the area in dimension $2$, we switch notation from $\mu_{i}$ to $a_{i}$ throughout this section. Its main purpose is to explain that in dimension $2$, the skeleton can be prescribed, which seems much harder in higher dimensions. We recall the notion of regular grid in a slightly more general setting.
@@
Let $(S,\omega)$ be a compact surface with boundary, and let $\Gamma$ be a regular grid that decomposes $S$ into $m$ discs of area $a_{i}$, in each of which we chose a point $p_{i}$. Then there exists a Liouville polarisation $\bigl(\{(p_{i},a_{i})\},\lambda\bigr)$ of $S$ whose skeleton is $\Gamma$. This Liouville form can be required to coincide with $(R-a_{i})d\theta$ in Darboux coordinates near each $p_{i}$.
In the case of a closed disc $\overline{D}(A)$ (which will be the example we will use later on) one may attempt to prove the proposition as follows: Just define a first tame Liouville form with residues $a_{i}$ at $p_{i}$ by $\lambda:=\alpha_{{\operatorname{st}}}-\sum a_{i}\,d\theta_{i}$, where $\theta_{i}$ is the pull-back by the translation of the vector $p_{i}$ of the standard angular coordinate in $\mathbb{C}\setminus\{0\}$. Then the periods of $\lambda$ vanish on each loop of $\Gamma$, so we can correct $\lambda$ by adding the differential of a smooth function in order to make $\lambda$ vanish on $\Gamma$. Then the Liouville flow of $\lambda$ fixes $\Gamma$ pointwise, and so $\bigl(\{(p_{i},a_{i})\},\lambda\bigr)$ provides a Liouville polarisation of $\overline{D}(A)$ whose skeleton contains $\Gamma$. Unfortunately, this proof does not constrain the skeleton enough: it could be strictly bigger than $\Gamma$. We therefore proceed in a different manner. Recall that $X_{\lambda}$ is tangent to $\ker\lambda$, since $\lambda(X_{\lambda})=d\lambda(X_{\lambda},X_{\lambda})=0$. Our strategy is therefore to first construct a (singular) foliation $\mathcal{F}$} to which we wish $X_{\lambda}$ to be tangent to. Once this “vanishing foliation” is constructed, the following lemma will readily provide the Liouville form itself.
@@
for the set of singularities of $\Gamma$ in the boundary of ${\mathcal{D}}_{i}$. For each $p_{i}$ we choose a disc $D(p_{i})$ centred at $p_{i}$ that is contained in ${\mathcal{D}}_{i}$. The translation $(D(p_{i}),p_{i})\to(D(\varepsilon_{p_{i}}),0)$ to the disc of the same area centred at the origin is a Darboux chart that induces coordinates $(R,\theta)$ on $D(p_{i})$. Similarly, by our assumption on $\Gamma$ we can consider for each $q\in\operatorname{Sing}(\Gamma)\setminus\partial D$ symplectic coordinates $\phi_{q}=(R,\theta)\colon D(q)\to D(\varepsilon_{q})$ on a disk around $q$ such that
@@
Also, when $q\in\operatorname{Sing}(\Gamma)\cap\partial D$, we have symplectic coordinates $(R,\theta)$ on a neighbourhood $D^{+}(q)$ of $q$ in $D$, with values in $\{R<\varepsilon_{q},\theta\in[0,\pi]\}$ such that
@@
The radial ray that cuts the sector $S_{i}(q)$ into two equal sectors is called a local separatrix at $q$ and is denoted $s_{i}(q)$. We choose the various (half) disks $D(p_{i})$, $D(q)$, $D^{+}(q)$ so small that they are mutually disjoint.
@@
For a point $q\in\operatorname{Sing}(\Gamma)\setminus\partial D$, let $\rho_{q}\colon D(\varepsilon_{q})\to\mathbb{C}$ be the “ramified covering”
@@
Note that $\rho_{q}$ is smooth and symplectic on $D(\varepsilon_{q})\setminus\{0\}$, but only continuous at 0. The sectors $S_{i}(q)$ around $q$ are the preimages under $\rho_{q}\circ\phi_{q}$ of $\{y\geq 0\}$ or $\{y\leq 0\}$, and the separatrices $s_{i}(q)$ are sent by $\rho_{q}\circ\phi_{q}$ to $\{x=0\}$. We define the foliation in $D(q)$ to be the pull-back by $\rho_{q}\circ\phi_{q}$ of the vertical foliation:
@@
[…]
@@
Figure 3.1. Notation, and the foliation on $D(\varepsilon_{q_{1}})$ for $m_{q_{1}}=3$.
@@
[…]
@@
Since ${\mathcal{F}}_{i}$ is a smooth foliation on ${\mathcal{D}}_{i}\setminus\bigl(\{p_{i}\}\cup{\mathcal{S}}_{i}\bigr)$ and by the smooth fitting of the leaves across the $\partial{\mathcal{D}}_{i}(q_{j}\to q_{j+1})$, the foliation $\mathcal{F}$} on $D\setminus\bigl(\bigcup\{p_{i}\}\cup\operatorname{Sing}(\Gamma)\bigr)$ defined by
@@
is $C^{\infty}$-smooth, and has a unique extension to a singular foliation on the whole disc $D$ whose singularities are $\bigcup\{p_{i}\}\cup\operatorname{Sing}(\Gamma)$. The different separatrices $s(p_{i}\to q)$, $q\in{\mathcal{S}}_{i}$, are smooth leaves. The leaves of $\mathcal{F}$} that pass through a point in $\Gamma\setminus\operatorname{Sing}(\Gamma)$ that separates between ${\mathcal{D}}_{i}$ and ${\mathcal{D}}_{j}$ join $p_{i}$ to $p_{j}$. Finally, the leaves through a sequence of regular points on ${\mathcal{D}}_{i}\cap{\mathcal{D}}_{j}$ that converge to a singular point $q\in\operatorname{Sing}(\Gamma)$ break into the two separatrices $s(p_{i}\to q)$ and $s(p_{j}\to q)$.
@@
By applying Lemma 3.3 twice, first to the curves of the foliation $\mathcal{F}$} of the interior $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$ that emanate from $p_{i}$ with $a:=a_{i}$, and then to the same set of curves but seen as emanating from $p_{j}$ with $a:=a_{j}$, we obtain on $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$ two Liouville forms $\lambda_{i},\lambda_{j}$. We claim that they coincide. To see this, we first check that these forms vanish on $\Gamma$. Indeed, let $q\in\Gamma\cap\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$. Since $\lambda_{i}$ vanishes along $\mathcal{F}$}, we already have $\lambda_{i}(q)v=0$ for $v\in T_{q}\gamma$, where $\gamma$ is the leaf through $q$. On the other hand, parametrise the leaves through $q\in\Gamma\cap\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$ by curves $\gamma(\theta,t)$, where $\theta$ represents the angle of the leaf at $p_{i}$ and the parameter $t$ verifies $\gamma_{\theta}(1)\in\Gamma$, and put $\frac{\partial}{\partial\theta}(q):=\frac{\partial\gamma}{\partial\theta}(\theta(q),1)\in T_{q}\Gamma$. Since $\lambda_{i}$ vanishes along $\mathcal{F}$}, and by Stokes’ theorem and the second property of $\lambda_{i}$ from Lemma 3.3,
@@
where ${\mathcal{A}}_{i}(\theta)$ is the area of the disc bounded by $\Gamma$, $\gamma(\theta,[0,1])$, and $\gamma(\theta(q),[0,1])$. By the choice of our foliation (point $5$ in Step 3), the right hand side of (3.1) vanishes, so $\lambda_{i}(q)$ vanishes. Now, since $\lambda_{i}$ and $\lambda_{j}$ coincide on $\mathcal{F}$} and on $\Gamma$, Stokes’ theorem shows that they coincide on $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$. We therefore define
$\lambda|_{\overset{\raisebox{-0.64583pt}[0.0pt][-0.64583pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})}\,:=\,\lambda_{i}=\lambda_{j}.$
Since $\lambda_{i}$ is defined and smooth on $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}_{i}\setminus\{p_{i}\}$, and not only on $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}(p_{i},p_{j})$, we can glue the $\lambda_{i}$ to get a smooth Liouville form $\lambda$ on $D\setminus\operatorname{Sing}(\Gamma)$. Notice that for $\gamma(\theta,t)$ as defined above, we have by the same argument
@@
Hence on $\overset{\raisebox{-0.90417pt}[0.0pt][-0.90417pt]{\mbox{$\scriptscriptstyle\circ$}}\mskip-2.0mu}{{\mathcal{D}}}_{i}\setminus\{p_{i}\}$ we have $d\lambda(X_{\lambda},\frac{\partial}{\partial\theta})=\lambda(\frac{\partial}{\partial\theta})<0$, which shows that the coordinate $t$ is (negative) gradient-like for the Liouville flow $X_{\lambda}$ of $\lambda$, and so ${\mathcal{B}}(p_{i},\lambda)\supset{\mathcal{D}}_{i}$. On the other hand, $\lambda$ vanishes on $\Gamma$, so $\Gamma$ is invariant under the Liouville flow of $\lambda$. Summarizing, we have therefore proved that
@@
    $\lambda$ is tame at each $p_{i}$ with residue $a_{i}$, in fact $\lambda=(R-a_{i})d\theta$ near $p_{i}$,
@@
[…]
@@
[…]
@@
The function $r^{m_{q}}\sin(2\pi m_{q}\theta)$ is smooth on $\mathbb{R}^{2}$, since it is the imaginary part of $z\mapsto z^{m_{q}}$. Let $\chi\colon[0,2\varepsilon_{q}]\to\mathbb{R}_{\geq 0}$ be a function that coincides with $r^{m_{q}}$ near 0 and with $R$ on $\varepsilon_{q},2\varepsilon_{q}]$, is positive and smooth except at 0, and meets $\chi(R)<R$ on $(0,\varepsilon_{q})$. Define the smooth 1-form $\lambda^{\prime}$ on $D$ by $\lambda^{\prime}=\lambda$ on $D\setminus\bigcup_{q\in\operatorname{Sing}\Gamma}D(q)$ and
@@
Then $\lambda^{\prime}$ is a smooth Liouville form on $D\setminus\{p_{1},\dots,p_{m}\}$, tame at $p_{i}$ with residue $a_{i}$. We claim that the skeleton of $\lambda^{\prime}$ is $\Gamma$, as required. To see this, notice that $X_{\lambda^{\prime}}=X_{\lambda}$ on $D\setminus\bigcup_{q\in\operatorname{Sing}\Gamma}D(q)$, and
@@
$X_{\lambda^{\prime}}$ is radial on $\Gamma\cap D(q)$ (hence its flow preserves $\Gamma$), and the trajectories of the points in $D(q)\setminus\Gamma$ flow away from $\Gamma$. From the $\frac{\partial}{\partial R}$-component, and from $\Gamma\supset\{\cos(2\pi m_{q}\theta)=1\}$ and $\chi(R)<R$ on $(0,\varepsilon_{q})$, we infer that the trajectories of the points in $D(q)\setminus\Gamma$ all leave $D(q)$, and hence reach the set $\{\lambda=\lambda^{\prime}\}\setminus\Gamma$ (since $\Gamma$ is invariant). It follows that $\textnormal{Skel}(\lambda^{\prime})=\Gamma$. $\square$
@@
Let $(S,\omega)$ be a compact symplectic surface with or without boundary. Let $\bigl({\mathcal{P}}:=\{(p_{i},a_{i}),\;i=1,\dots,n\},\lambda\bigr)$ with $p_{i}\in S\setminus\partial S$ be a polarisation of $S$ (where $\lambda$ is a tame Liouville form with residues $a_{i}$ at $p_{i}$ and with flow tangent to $\partial S$), and set $\Gamma:=\textnormal{Skel}(S,{\mathcal{P}},\lambda)$.
@@
Then $\lambda^{\prime}$ is a Liouville form tame at $p_{i}^{j}$ with residue $a_{i}^{j}$. Since it coincides with $\lambda$ on $\Gamma\cup\partial S$, its flow preserves both $\Gamma$ and $\partial S$, so $\left({\mathcal{P}}^{\prime}:=\left\{(p_{i}^{j},a_{i}^{j})\right\},\lambda^{\prime}\right)$ is a Liouville polarisation of $S$. Since $\lambda^{\prime}$ coincides with $\lambda$ outside the discs $D_{i}$,